OpenStudy (anonymous):

How old will Bill be in 10 years if he is x + 2 years old now? Which expression represents this? (x + 2) - 10 (x + 2) + 10 10(x + 2)

OpenStudy (freckles):

pretend you are 6 six year old right now how old will you be in 10 years?

OpenStudy (freckles):

if you are 6 years old right now then next year you would be 6+1=7 if you are 6 years old right now then in two years you would be 6+2=8 if you are 6 years old right now then in three years you would be 6+3=9 if you are 6 years old right now then in n years you would be 6+n -- so if you are (x+2) years old right now then in 10 years you would be ?

OpenStudy (bee_see):